Acetic acid is a weak acid (HC₂H₃O₂). It is the active ingredient in vinegar. If the initial concentration of acetic acid is 0.18 M and the equilibrium concentration of the hydronium ion is 0.0021 M, calculate Ka for acetic acid.
1
Expert's answer
2020-05-10T14:57:42-0400

"Ka= \frac{[H3O6^+].[CH3COO^-]}{[CH3COOH]}"

Ka=0.0021/0.18

Ka=0.0117

Ka=1.7*10-3